OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 2) and is parallel to y = 4x + 7. y = 1 over 4 x + 5 over 2 y = -4x - 6 y = 4x + 10 y = -x + -1 over 4 + 3 over 2 I don't get math.

OpenStudy (ajprincess):

What is the slope of the equation y=4x+7? Can u plz tell me.

OpenStudy (anonymous):

m=4

OpenStudy (ajprincess):

For parallel lines slope is the same. So the slope of the required equation is also 4. Also \(\large\frac{y-y_1}{x-x_1}=m\) That is \(\large\frac{y-2}{x-(-2)}=4\) Nw can u find the equation?

OpenStudy (anonymous):

since it is parallel to y = 4x + 7.its slope is m=4 since slope of parallel lines are equal slope of required line must be m=4 now using the point slope form \[\Large y-y_{0}=m(x-x_{0})\] so \[\Large y-2=4(x-(-2))\] can you now make it as point slope form just simplify and take the -2 from the left side to the right.

OpenStudy (anonymous):

@pam142 can you find its slope intercept form now?

OpenStudy (anonymous):

so I add 2?

OpenStudy (anonymous):

ok let me get you one step ahead . \[\Large y-2=4(x+2)\] \[\Large y-2=4x+8\] can you make the following form from the above \[\Large y=mx+c\] ??

OpenStudy (anonymous):

y=4x+6?

OpenStudy (anonymous):

yeah just add 2 to the both sides and -2 will get cancelled on the left side and added to the right and you will have y=4x+10 !

OpenStudy (anonymous):

fat baby

OpenStudy (anonymous):

ohhh okay thanks :D

OpenStudy (anonymous):

you 're welcome:)